OpenStudy (anonymous):

if a principal, P, is invested at r% interest compounded annually then its future value, S, after n years is given by \(\large S=P(1+ \frac{r}{100})^n \) Use this formula to show that if an interest rate of r% is compounded k times a year, then after t years \(\large S=P(1+ \frac{r}{100k})^kt \) b) show that if m=100k/r then the formula in part (1) can be written as S+P((1+1/m)^m)^rt/100 c) use the difinition e=lim m>00(1+1/m)^m to deduce that if the interest is compounded with every- increasing frequency ( that is conitnuously ) then s=Pe^rt/100 oo=infinity thank you for hellping me

OpenStudy (anonymous):

i need some help ^^

OpenStudy (anonymous):

i am not sure what it means by "show" but if your interest rate is \(r\) then if it is compounded \(k\) times per year, each \(k\)th part of a year you get \(\frac{r}{k}\) interest. then if the number of years is \(t\) there are \(kt\) compounding periods, giving you the formula you wrote above

OpenStudy (anonymous):

i said let n= t and k=1 so anwser gona be s=p(1+r)^n

OpenStudy (anonymous):

i am not sure too that why i am asking

OpenStudy (anonymous):

the second part is a replacement \(\large S=P(1+ \frac{r}{100k})^{kt} \) put \(m=\frac{100k}{r}\) which makes \(\frac{1}{m}=\frac{r}{100k}\)

OpenStudy (anonymous):

this 100 is annoying, \(r\) should be written as a number, not a percent. but if that is what is says, that is what you have to do

OpenStudy (anonymous):

yea it like that written i know it is annoying

OpenStudy (anonymous):

if you replace \(\frac{r}{100k}\) by \(\frac{1}{m}\) you get \(\large S=P(1+ \frac{1}{m})^{kt} \)

OpenStudy (anonymous):

so it gona be the final anwser S=p(1+1/m)^kt after replacing m=100k/r right?

OpenStudy (anonymous):

that was wrong, sorry

OpenStudy (anonymous):

kk how i suppose submit in my paper how i start to writte i understand what you say how i should writte in my paper^^

OpenStudy (anonymous):

since \[m=\frac{100k}{r}\] you get \[k=\frac{mr}{100}\]

OpenStudy (anonymous):

to your exponent is \[{kt}=\frac{mrt}{100}=m\times \frac{rt}{100}\]

OpenStudy (anonymous):

this gives \[S+P((1+1/m)^m)^{rt/100}\]

OpenStudy (anonymous):

with an equal sign not a plus sign

OpenStudy (anonymous):

then as the \(k\to \infty\) you have \(m\to \infty\) i.e. as the compounding period gets shorter and shorter, i.e. as the number of compounding periods gets larger and larger, you get \[\lim_{m\to \infty}P(1+\frac{1}{m})^m)^{rt/100}\]\[=Pe^{rt/100}\]

OpenStudy (anonymous):

i would write the reasoning in english and the algebra as algebra

OpenStudy (anonymous):

that the anwser for c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, b and c

OpenStudy (anonymous):

b is algebra, c is taking the limit to get \(e\)

OpenStudy (anonymous):

for b i write like that since m=100kr you get k=mr100 kt=mrt100=m×rt100 S+P((1+1/m)m)rt/100

OpenStudy (anonymous):

for c ) like that then as the k→∞ you have m→∞ i.e. as the compounding period gets shorter and shorter, i.e. as the number of compounding periods gets larger and larger, you get limm→∞P(1+1m)m)rt/100 =Pert/100

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

i should righ like that satellite?